find the approximate change in the volume of 1/(x)2 .when x changes from x=2 to x= 2.002

Grade:12

1 Answers

Arun
6 years ago
Dear Abhishek
 
f(x) = 1/x²
f'(x) = -2/x³
 
Approximate change in y = f'(x) * \Delta x
 = -2/2³  * (0.002)
 = -0.002/4
 = -0.0005
 
 
Regards
Arun (askIITians forum expert)